The transverse displacement (y) of a wave is given as a function of position (x in meters) and time (t in seconds) by the expression to the right. Determine the wavelength, frequency, period, and phase constant of this waveform.

y(x,t)= y_{m} sin(0.333x + 3.38 + 801t)
Answer #1

Mas per it rev L 1.9チ 655A/m JF Fourth harmonic= n/d L 44) 2 rep. 355 = 36 Hz. 0.985